OpenStudy (anonymous):

How do I use partial fractions to find the integral of 1/(x^2 - x) dx?

OpenStudy (anonymous):

I have the integral of 1/ (x-1)(x) = A/(x-1) + B/(x) So the integral of 1 = A(x) + B(x+1)? I don't know if that's right, and I don't know what to do next.

OpenStudy (phi):

you have \[ \frac{1}{x(x-1)}= \frac{A}{(x-1)}+ \frac{B}{x} \] or \[ \frac{1}{x(x-1)}= \frac{Ax}{x(x-1) }+\frac{B(x-1)}{x(x-1)}\] or \[ \frac{1}{x(x-1)}= \frac{Ax+B(x-1)}{x(x-1) }\] looking at this, we see that the tops of both sides must be equal \[ 1= Ax+B(x-1) \] or \[ (A+B)x -B = 1 \] or \[ (A+B)x -B = 0x+ 1 \] now match corresponding coefficients to find 2 equations. in other words, the coefficient on the x terms must be equal: A+B= 0 and the constant terms must be equal -B= 1 now solve for A and B. use them in your original fractions which can be integrated (you get a natural log)

OpenStudy (anonymous):

Cool! I was just missing the rearranging step from (x-1)B+Ax=1 to (A+B)x-B=1. Now it makes sense. Thanks! :D